Question 3 Rossi Incorporated makes track suits that sell for coming year. 70 each. Actual sales are $1,036,000. Management estimates that fixed costs will total $233,100 and variable costs will be $49 per unit this Calculate the break even point n sales dollars using the contribution margin ratio places, e.g. 125.) Round contribution margin ratio to 4 decimal places, eg. 15.2964% and final answer to 0 decimal Break-even point in dollars LINK TO TEXT Calculate the margin of safety in dollars and the margin of safety ratio. (Round margin of safety in dollars to the nearest whole dollar, e.g. 5,275 and margin of safety ratio to 2 decimal places, c.g. 15.25%.) Margin of safety Margin of safety ratio LINK TO TEXT Calculate the sales dollars required to earn an operating income of $184,500. (Round answer to the nearest whole dollar, e.g. 5,215.) Required sales in dollars $ LINK TO TEXT
Answer #1

Contribution margin ratio = (70-49)/70 = 30%

a) Break even point in dollars = Fixed cost / Contribution margin ratio

= 233100/0.30

Break even point in dollars = $777000

b) Margin of safety = Actual Sales - Break even sales

= 1036000-777000

Margin of safety = $259000

Margin of safety ratio = 259000/1036000 = 25%

c) Required sales in dollars = (233100+184500)/0.30 = $1392000